OpenStudy (anonymous):

Find a third degree polynomial function of the lowest degree that has the zeros below and whose leading coefficient is on. 5,-7-√2,-7+√2

OpenStudy (mertsj):

\[(x-5)[x-(-7-\sqrt{2})][x-(-7+\sqrt{2})]=0\]